# CoT-UQ: Improving Response-wise Uncertainty Quantification in LLMs with Chain-of-Thought

### v1

@@ 4.2 Stage 2 / Step 4: Reasoning Enhanced Uncertainty Quantification Strategy
Table 1: AUROC ($\uparrow$) comparison of Llama 3.1-8B and Llama 2-13B on various benchmarks for logical and mathematical reasoning tasks, where AP indicates aggregated probabilities and SE denotes self-evaluation.
@@ 5.1 Experimental Setups
We consider the following reasoning scenarios: logical reasoning and mathematical reasoning, where existing UQ strategies suffer from overconfidence issues. Specifically, for logical reasoning, we use the HotpotQA dataset (Yang et al., 2018) and the 2WikiMultiHopQA dataset (Ho et al., 2020); for mathematical reasoning, we use the GSM8K dataset (Cobbe et al., 2021), the SVAMP dataset (Patel et al., 2021), and the ASDiv dataset (Miao et al., 2021). For models, we use Llama2-13B and Llama3.1-8B (Touvron et al., 2023). Details of dataset statistics are provided in Appendix A.1, while the implementation details of our method, including LLMs’ hyper-parameters and prompts are presented in Appendix B.
@@
Figure 4: Comparison of different implementations of CoT-UQ based on Self-Probing. KEYKeywords consistently achieves better performance on logical reasoning tasks HotpotQA and 2WikiMultiHopQA), while KEYStep performs better in mathematical reasoning (GSM8K and ASDiv).
@@ 5.2 Main Results
We report the results of recommended realizations in Table 1 and a detailed explanation and analysis for the above suggestions in Section 5.3. The results of these five datasets demonstrate an average of +4.4% improvement, highlighting the effectiveness of CoT-UQ compared to standard SE baselines. Notably, CoT-UQ applied to SE strategies shows a greater performance improvement in mathematical reasoning tasks (+5.3%) compared to logical reasoning tasks (+3.5%). This suggests that incorporating reasoning in UQ allows the model to identify critical errors in the thought process during self-evaluation, especially in mathematical problems where a single misstep can lead to incorrect conclusions. We provide detailed evidence and analysis via case studies in Appendix C.2.
@@ 5.3 Ablation and Future Discussions
In this section, we provide a thorough understanding of our CoT-UQ. Additional results and discussions (e.g., sensitivity to hyper-parameters and case studies) can be found in Appendix C.
@@ Appendix B Implementation Details
Prompts in Stage 2 of CoT-UQ.
  • ALLSteps:
  • KEYStep:
[…]
B.3 Computational Costs Analysis
Appendix C Additional Experimental Results and Further Discussion
C.1 Sensitivity to Keywords Filtering threshold in SE.
Figure 6: Analysis of the sensitivity to the importance filtering threshold $\tau$ on the HotpotQA benchmark. The experiments are based on LLaMA3-8B.

### v2

@@ 4.2 Stage 2 / Step 4: Reasoning Enhanced Uncertainty Quantification Strategy
Table 1: AUROC ($\uparrow$) comparison of Llama 3.1-8B and Llama 2-13B on various benchmarks for logical and mathematical reasoning tasks, where AP indicates aggregated probabilities and SE denotes self-evaluation. For the chosen type of SE strategy, we propose KEYKeywords for both two SE strategies in logical reasoning tasks, ALLSteps for P(True) and KEYStep for Self-Probing in mathematical reasoning tasks (See Section 5.2 for analysis).
@@ 5.1 Experimental Setups
We consider the following reasoning scenarios: logical reasoning and mathematical reasoning, where existing UQ strategies suffer from overconfidence issues. Specifically, for logical reasoning, we use the HotpotQA dataset (Yang et al., 2018) and the 2WikiMultiHopQA dataset (Ho et al., 2020); for mathematical reasoning, we use the GSM8K dataset (Cobbe et al., 2021), the SVAMP dataset (Patel et al., 2021), and the ASDiv dataset (Miao et al., 2021). For models, we use Llama 2-13B and Llama 3.1-8B (Touvron et al., 2023). Details of dataset statistics are provided in Appendix A.1, while the implementation details including LLMs’ hyper-parameters, prompts, and computational costs are provided in Appendix B.
@@
Figure 4: Comparison of different implementations of CoT-UQ based on Self-Probing. KEYKeywords consistently achieves better performance on logical reasoning tasks (HotpotQA and 2WikiMultiHopQA), while KEYStep performs better in mathematical reasoning (GSM8K and ASDiv).
@@ 5.2 Main Results
We report the results of recommended realizations in Table 1 and a detailed explanation and analysis for the above suggestions in Section 5.3. The results of these five datasets demonstrate an average of +4.4% improvement, highlighting the effectiveness of CoT-UQ compared to standard SE baselines. Notably, CoT-UQ applied to SE strategies shows a greater performance improvement in mathematical reasoning tasks (+5.3%) compared to logical reasoning tasks (+3.5%). This suggests that incorporating reasoning in UQ allows the model to identify critical errors in the thought process during self-evaluation, especially in mathematical problems where a single misstep can lead to incorrect conclusions. We provide detailed evidence and analysis via various case studies in Appendix C.4.
@@ 5.3 Ablation and Future Discussions
In this section, we provide a thorough understanding of our CoT-UQ. Additional results and discussions (e.g., comprehensive validation across model families and metric, sensitivity to hyper-parameters, importance scores visualization, and several case studies) can be found in Appendix C.
@@ Appendix B Implementation Details
Prompts in Stage 2 of CoT-UQ.
  • ALLSteps:
  • KEYStep:
[…]
B.3 Computational Costs Analysis
Additionally, we further analyze the actual computational overhead associated with CoT-UQ compared to the baseline methods. The baselines in our experiments are essentially not “single-pass”. They inherently involve two inference steps: answer generation and uncertainty aggregation/self-evaluation. The additional overhead introduced by CoT-UQ only comes from reasoning chain generation and keyword extraction/scoring. Based on our empirical setting, generating an answer alone requires 3 (± 0.2) seconds per question, incorporating the CoT adds 3.5 (± 0.3) seconds, and the subsequent keyword extraction and scoring steps add 3.5 (± 0.5) seconds. Overall, CoT-UQ incurs less than 8 seconds of additional computational overhead per question. In contrast, sampling-based UQ methods generally require at least 10 answers per question to ensure robust performance Qiu and Miikkulainen (2024), leading to over 30 seconds per question. Thus, our method is significantly more efficient in practice.
Appendix C Additional Experimental Results and Further Discussion
C.1 Comprehensive Validation Across Model Families and Metrics
In this part, we present experimental results to further validate the generalizability, effectiveness, and compatibility of our approach.
To validate the generalizability of our approach, we conduct experiments based on different model families since uncertainty quantification performance can vary significantly across model families. Specifically, Table 3 explicitly reports the performance of CoT-UQ based on Mistral 7B across various tasks and datasets compared to baseline methods. The results provide further evidence of the general applicability of our proposed approach.
In addtion, to comprehensively demonstrate the validity of our method, we report experimental results including the AUPRC metric in Table 5. These results indicate that our CoT-UQ approach consistently outperforms baseline methods under AUPRC evaluation, further confirming its effectiveness and robustness.
Furthermore, we investigate whether incorporating CoT for fact-checking affects the overall performance of the models. As shown in Table 5, our findings reveal that CoT in fact provides consistent accuracy improvements across different datasets. It is also important to note that all UQ methods in the paper, including CoT-UQ, do not affect the model’s accuracy, as the uncertainty indicator is computed/evaluated after the model has generated its response.
Strategy Method HotpotQA 2WikiMHQA SVAMP ASDiv
AP Probas-mean 54.63 58.19 59.88 57.26
w/ CoT-UQ 58.24 65.58 65.83 67.40
TOKENSAR 55.62 55.60 60.69 58.43
w/ CoT-UQ 58.59 66.06 65.42 68.18
SE P(True) 48.02 47.13 42.93 48.44
w/ CoT-UQ 50.18 49.78 50.37 51.30
Table 3: AUROC ($\uparrow$) Performance Comparison with baselines based on Mistral 7B.
Method HotpotQA ASDiv
AUROC AUPRC AUROC AUPRC
Probas-mean 53.73 29.14 58.34 69.03
w/ CoT-UQ 62.01 32.25 64.52 73.95
TOKENSAR 53.57 28.41 58.71 68.60
w/ CoT-UQ 61.07 31.27 66.91 75.43
P(True) 62.39 35.68 47.23 60.28
w/ CoT-UQ 63.10 37.29 53.20 64.15
Table 4: Comparison with baselines including AUPRC ($\uparrow$) evaluation based on Llama 3.1-8B.
Datasets w/o CoT w/ CoT
Acc. AUROC Acc. AUROC
HotpotQA 34.2 56.27 37.1 66.56
2WikiMHQA 28.2 51.54 32.4 63.29
GSM8K 17.8 53.96 18.9 58.54
SVAMP 38.6 54.48 42.2 57.37
ASDiv 42.4 57.73 47.8 59.44
Table 5: Accuracy ($\uparrow$) w/ and w/o CoT based on Llama 2-13B. We report the AUROC results based on Probas-mean method.
C.2 Sensitivity to Keywords Filtering threshold in SE
Figure 6: Analysis of the sensitivity to the importance filtering threshold $\tau$ on the HotpotQA benchmark. The experiments are based on Llama 3.1-8B.
Figure 7: Frequency distribution of importance scores assigned by Llama 3.1-8B on the HotpotQA dataset.
C.3 Visualization of Model’s Importance Scores
We presented the frequency distribution of importance scores assigned by Llama 3.1-8B to keywords extracted from reasoning chains in the HotpotQA dataset, as shown in Figure 7. The results show that the LLM assigns diverse importance scores to the extracted keywords. The distribution is not concentrated on a single score but spans across the entire spectrum, indicating that the model is capable of differentiating the relative importance of different keywords. This demonstrates that LLMs are able to evaluate which parts of the reasoning process are more critical for the final answer.